OpenStudy (anonymous):

How do I find the x and y intercepts of x-2y=4?

OpenStudy (anonymous):

To find an intercept, set one of the variables equal to zero. When x = 0, 0 - 2y = 4 -2y = 4 -y = 2 y = -2 (0, -2) is the y - intercept. Use the same process, but with y = 0 for the x - intercept.

OpenStudy (anonymous):

Thanks much!